OpenStudy (dls):

What is the principal solution for tanx=sqrt{3}?

OpenStudy (dls):

\[\LARGE x=\tan^{-1} 3\] And range for tan inverse is -pi/2 to +pi/2 so it should be pi/3?

OpenStudy (dls):

But the answer is pi/3 and 4pi/3 but 4pi/3 is not in the range for tan inverse function..why is it so?

OpenStudy (dls):

I researched a bit,and found principal solution takes values from o to 2pi,but that is quite contradicting

terenzreignz (terenzreignz):

I think principal means within the interval \[\large [0 \ , 2\pi)\]

terenzreignz (terenzreignz):

Well, think about it, both are solutions. And anything bigger than 2pi is coterminal to one of them.

OpenStudy (dls):

yeah,but then x=tan inverse root 3 and tanx=sqrt{3} have different principal solutions?that is the main confusing question to me.

OpenStudy (dls):

both seem same things

terenzreignz (terenzreignz):

Well, for one thing... the solution was for this equation... \[\large \tan \ x = \sqrt3\]and not \[\large x = \tan^{-1}\sqrt{3}\] It's jut like how this \[\large x^2 = 4\]has two solutions but \[\large x = \sqrt4\] has only one.

OpenStudy (dls):

but still both are the same things? :/

terenzreignz (terenzreignz):

They're not. Because to much like the square root, the inverse tangent only becomes a function through a restricted domain.

OpenStudy (dls):

what did this mean^^?

OpenStudy (dls):

don't play with big words :|

terenzreignz (terenzreignz):

It takes a deep understanding of functions :D Take the squaring function... \[\huge f:\mathbb{R}\rightarrow\mathbb{R}\]\[\huge f:x\rightarrow x^2\]

terenzreignz (terenzreignz):

Domain is all real numbers right?

OpenStudy (dls):

yes!

terenzreignz (terenzreignz):

Now, the function becomes onto, if we restrict the range... \[\huge f:\mathbb{R}\rightarrow [0,\infty)\]

terenzreignz (terenzreignz):

We need this function to be onto (surjective) if we're to define an inverse relation.

OpenStudy (dls):

I see! cleared a lot of things! thanks :)

OpenStudy (dls):

@yrelhan4,this will clear your doubt too :)

terenzreignz (terenzreignz):

I wasn't even finished yet :/ But lol, you seem to understand it now :D

OpenStudy (dls):

yeah,:P